OpenStudy (anonymous):

If a sample of pure silver (Ag) contains 3.35 x 10^22 atoms of iodine, how many moles of silver are in the sample? a. 17.9 mol b. 2.82 mol c. 0.180 mol d. 0.0556 mol

OpenStudy (anonymous):

\[\frac{3.35x10^{22}}{6.022x10^{23}}=\]

OpenStudy (lgbasallote):

if you'd like to see the dimension analysis itself... \[\large 3.35 \times 10^{22} \cancel{\text{atoms of iodine}} \; \times \; \frac{1 \text{mole of iodine}}{6.022 \times 10^{23} \cancel{\text{atoms of iodine}}}\]

OpenStudy (lgbasallote):

change that 1 mole of iodine into 1 mole of silver

OpenStudy (anonymous):

D 00556

OpenStudy (anonymous):

boo